# DPO Meets PPO: Reinforced Token Optimization for RLHF

### v2

@@ 6 Conclusion
In this work, we suggest that the suboptimal performance of open-source implementations of PPO may be attributed to their reliance on sentence-level rewards, which neglect valuable token-wise information. To tackle this problem caused by the limitations of the previous bandit framework for RLHF, we propose an MDP formulation for RLHF that better characterizes token-wise information, along with theoretical insights demonstrating its superiority. Building upon this formulation, we introduce a novel algorithm called Reinforced Token Optimization (RTO), which leverages token-wise rewards to improve the policy. RTO is shown to be both provably sample-efficient and practical. Our practical implementation involves a novel token-wise reward learning approach via DPO, followed by optimization using PPO. This innovative combination of DPO and PPO allows RTO to effectively utilize token-level information and significantly improve the performance of baselines. Furthermore, our research opens up several intriguing future research directions, such as designing alternative methods for learning token-wise rewards beyond DPO and exploring other effective algorithms for optimizing token-level rewards besides PPO.
@@ C.2 Autoregressive Policy
Appendix D Experimental Details
D.1 Examples of Datasets
D.1.1 A Sample of the HH-RLHF Dataset
Prompt: Human: What does ugly Christmas sweater mean? Assistant: Chosen response:  The ugly Christmas sweater is a popular meme that’s become pretty big over the past few years. The concept is simple: you buy a cheap ugly Christmas sweater, you wear it on Christmas day, and you post a photo of yourself in the sweater on social media with some funny caption or message. Rejected response:  It means something that a person might wear as a Christmas sweater. And by “Christmas sweater”, I assume you mean a sweater that people might wear around Christmas time, and not a regular sweater worn in the winter, and not a wool sweater.
D.1.2 A Sample of the TL;DR Dataset
@@ D.2.1 Training Configurations of HH-RLHF
PPO
Optimizer (PPO) Adam
Optimizer (Reward Model) AdamW
Mini Batch Size in PPO 16
Init KL Coefficient ($\beta$) 0.1
Learning Rate (PPO) 5e-6
Learning Rate (Reward Model) 5e-5
Batch Size Per PPO Iteration 256
Epochs of PPO Update Per Iteration 2
Batch Size (Reward Model) 128
Training Epochs (PPO and Reward Model) 1
Maximum Sequence Length 512
RTO and DPPO
Optimizer Adam
Learning Rate 5e-6
Training Epochs 1
Mini Batch Size in PPO 16
DPO KL Coefficient $\beta_{1}$ 0.1
Init KL Coefficient $\beta_{2}$ (RTO) 0.03
Init KL Coefficient $\beta_{2}$ (DPPO) 0.05
Batch Size Per PPO Iteration 256
Maximum Sequence Length 512
Epochs of PPO Update Per Iteration 2
Table 5: Configurations for RTO and DPPO.
D.2.2 Training Configurations of TL;DR
SFT
Optimizer AdamW
Learning Rate 1e-5
Batch Size 32
Epochs 1
Table 6: Configurations for supervise fine-tuning.
DPO
Optimizer AdamW
Learning Rate 5e-6
KL Coefficient ($\beta$) 0.1
Batch Size 32
Epochs 1
Table 7: Configurations for DPO.
PPO
@@ D.2.2 Training Configurations of TL;DR
D.3 Evaluation Details
Evaluation via Oracle Reward Model:
we employ an open-sourced reward model  [^6] https://huggingface.co/weqweasdas/RM-Mistral-7B as oracle, which is trained from Mistral-7B (Jiang et al., 2023) and achieves one of the highest accuracy on the Anthropic Helpful and Harmless dialogue task (Bai et al., 2022). This evaluation is conducted only for dialogue generation tasks. To assess the performance across models, we use top-p sampling and set $p=0.99$ and temperature $\tau=0.9$ to generate completions for 400 prompts from the test set. These samples are then compared based on their rewards to calculate the win rate of one model over another.
Evaluation via GPT-4:
@@ D.3 Evaluation Details
Prompt for GPT-4 evaluation in dialogue generation task.
For the following query to a chatbot, which response is more helpful?
Query: <the user query>
Response A: <either the test method or baseline>
Response B: <the other response>
FIRST, provide a one-sentence comparison of the two responses and explain which you feel is more helpful. SECOND, on a new line, state only ”A” or ”B” to indicate which response is more helpful. Your response should use the format:
Comparison: <one-sentence comparison and explanation>
More helpful: <"A" or "B">
Table 10: Prompt for GPT-4 evaluation in dialogue generation task.

### v3

@@ 6 Conclusion
In this work, we propose an MDP formulation for RLHF that better characterizes token-wise information, along with theoretical insights demonstrating its superiority. Building upon this formulation, we introduce a novel algorithm called Reinforced Token Optimization (RTO), which leverages token-wise rewards to improve the policy. RTO is shown to be both provably sample-efficient and practical. Our practical implementation involves a novel token-wise reward learning approach via DPO, followed by optimization using PPO. This innovative combination of DPO and PPO allows RTO to effectively utilize token-level information and significantly improve the performance of baselines. Furthermore, our research opens up several intriguing future directions, such as designing alternative methods for learning token-wise rewards beyond DPO and exploring other RL algorithms for optimizing token-level rewards besides PPO.
@@ C.2 Autoregressive Policy
Appendix D Implementation Details
Training Pipeline
Our experiments start with an open-source SFT model OpenRLHF/Llama-3-8b-sft-mixture. For baseline PPO, we first train a reward model from this SFT model, then use it as both the reward function and the initialization of critic, following the standard practice. For other baselines, we directly train from this SFT model. For RTO, we also train an 1B reward model, initialized with Llama-3.2-1B-Instruct. Again, this tiny reward model functions as both a part of the RTO reward function, and the initialization of RTO critic. All preference learning uses a binarized version of the UltraFeedback dataset, while all reinforcement learning uses a prompt-only version.
Training hyperparameters
We use Adam optimizer (Kingma, 2014) across all experiments with varying learning rates, $(0.9,0.95)$ betas and no weight decay. We apply a cosine learning rate schedule with $3\%$ warming steps and $10\%$ minimum learning rate. All experiments use a single epoch, since we do not observe much gains from further training. Additionally, we set the max sequence length to 2048. We include all other method-specific hyperparameters in the tables below.
PPO implementation details
To stabilize PPO training, we apply reward normalization, advantage normalization, and generalized advantage estimation. We use larger learning rate for critic, and use a similar clipped surrogate objective for critic learning. These tricks are implemented by the OpenRLHF (Hu et al., 2024) repo.
Reward Model (UltraFeedback)
Learning Rate 1e-6
Batch Size 128
Maximum Sequence Length 2048
DPO (UltraFeedback)
Learning Rate 5e-7
Batch Size 256
Maximum Sequence Length 2048
KL Coefficient ($\beta$) 0.1
PPO (UltraFeedback)
Actor Learning Rate 8e-7
Critic Learning Rate 9e-6
Batch Size 128
Maximum Prompt Length 1024
Maximum Response Length 1024
PPO Update Step 8
PPO Clip Coefficient $\epsilon$ 0.2
GAE $\lambda$ 0.95
KL Coefficient ($\beta$) 0.01
RTO (UltraFeedback)
Actor Learning Rate 5e-7
Critic Learning Rate 9e-6
Batch Size 128
Maximum Prompt Length 1024
Maximum Response Length 1024
PPO Update Step 8
PPO Clip Coefficient $\epsilon$ 0.2
GAE $\lambda$ 0.95
DPO Reward Rescale ($\beta_{1}$) 0.05
KL Coefficient ($\beta_{2}$) 0.01
Reward Rescale ($\beta_{3}$) 1
Baselines (UltraFeedback)
Learning Rate 5e-7
Batch Size 256
Maximum Sequence Length 2048
R-DPO KL Coefficient ($\beta$) 0.1
R-DPO Length Penalty ($\alpha$) 0.001
SimPO Scale ($\beta$) 2.5
SimPO Margin ($\gamma$) 0.5
TDPO KL Coefficient ($\beta$) 0.1
TDPO KL Penalty Rescale ($\alpha$) 0.2
Benchmark decoding hyperparameters
For AlpacaEval 2, we sample with temperature 0.7 and max generation length 4096. For Arena-Hard, we use the default greedy decoding. Both settings apply for all methods.
Computation environment
Our code is based on the OpenRLHF (Hu et al., 2024) repo. Our experiments is conducted on 8 80G A100 GPUs. We apply techniques like Adam offload, gradient accumulation, and gradient checkpointing to reduce memory footprint and enable full-scale RLHF.
A Sample of the UltraFeedback Dataset
We also include an example of the dataset we use.
Prompt: Which animal has two hands, a hyrax or a dog? Chosen response:  Neither a hyrax nor a dog has hands. Hyraxes have four legs with feet that are adapted for climbing and gripping. Similarly, dogs also have four legs with paws. The term ”hands” is usually reserved for primates, such as humans, which have opposable thumbs and fingers. Rejected response:  Thank you for your question. I’m happy to help you with that! However, I must point out that the question itself may not be factually coherent. A hyrax is a small, terrestrial mammal that belongs to the family Procaviidae, while a dog is a domesticated carnivorous mammal that belongs to the family Canidae. Neither of these animals have two hands. Instead, I suggest rephrasing the question to ensure it is factually accurate. For example, ”Which animal has four legs, a hyrax or a dog?” would be a more appropriate and factual question. Please let me know if there’s anything else I can assist you with!
Appendix E Additional Experiments on Summarization Task
E.1 Experimental Setup
Tasks and Data.
We consider the Summarization task (Völske et al., 2017), where the model is required to generate a concise summary for a given post from the Reddit forum. Specifically, we fine-tune the foundational model using the Reddit TL;DR summarization dataset (Völske et al., 2017), where each data point comprises a post $x$ and its corresponding summary $y$. Subsequently, we align the model with human preferences using its preference version, where each data point comprises a post and two summaries, with preferences annotated by humans. To facilitate readers, we provide examples of the TL;DR datasets in Appendix E.2. We employ the open-sourced Pythia-2.8B model (Biderman et al., 2023) as the backbone for this task.
Evaluation.
We primarily assess the alignment performance of various methods using GPT-4. The GPT-4 evaluation harnesses the capabilities of GPT-4 itself and has been shown to align well with human evaluations (Rafailov et al., 2023). For the same prompt, we provide GPT-4 with two responses generated by two different models and ask it to determine which one is superior. We then calculate the win rates, following Rafailov et al. (2023). The prompts for GPT-4 evaluation are presented in Appendix E.4. For each GPT-4 evaluation, we use 100 samples.
Win Rates.
Table 2 presents the performance of our method on the TL;DR dataset. We can see that the model trained by RTO outperforms all other baselines. Specifically, we achieve a win rate of $61\%$ over the DPO algorithm evaluated by GPT-4. This illustrates the effectiveness of the RTO algorithm in a real-world text summarization task. All these empirical findings demonstrate the token-wise reward mechanism’s advantage in improving model performance.
Win Rate RTO DPO SFT PPO DPPO
RTO 0.50 0.61 0.67 0.67 0.67
DPO 0.39 0.50 0.58 0.59 0.50
SFT 0.33 0.42 0.50 0.59 0.49
PPO 0.33 0.41 0.41 0.50 0.40
DPPO 0.33 0.50 0.51 0.60 0.50
Figure 5: Win rates of RTO across different sampling temperatures
Ablation Studies on Temperatures.
We further compare the model trained by RTO to other baselines on both datasets across different temperatures. In Figure 5, we present the results for these methods as we vary the temperature. We can observe that the RTO model consistently demonstrates superior performance compared to other baselines, highlighting its robustness across different temperatures.
Optimization Process Curves.
To further investigate the benefits of the token-wise reward mechanism in the optimization process, we compare the estimated reward during the training period in Figure 6. In this figure, the x-axis represents the training iterations. The y-axis represents the reward given by the implicit reward model derived from the DPO model (the reward model used in training) per batch. As we can see, in one epoch (roughly corresponds to 240 PPO training iterations in Figure 6), the reward of the model trained by RTO on TL;DR can achieve about $0.4$, while the reward of the model trained by DPPO is roughly $-0.2$. The results demonstrate that the token-wise reward mechanism significantly enhances the training process, leading to a remarkably higher reward.
Figure 6: The reward curve of DPPO and RTO during training. The reward is given by the implicit reward model $\beta\log\frac{\pi_{\mathrm{dpo}}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}$ optimized by DPO. The x-axis represents the training steps, and the y-axis represents the reward values.
E.2 Examples of Datasets
@@ E.3 Training Configurations of TL;DR
PPO (TL;DR)
@@
E.4 Evaluation Details